please help with this problem.
Solve the equation.
x^2 - x = 90

Answer by hamda(7)

x^2 - x = 90
x^2 - x - 90 = 0
x^2 - 10x + 9x - 90 = 0
x(x - 10) + 9(x - 10) = 0
(x - 10)(x + 9) = 0
x = 10 , x = -9 (Answer)